OpenStudy (anonymous):

Can someone help me find the range of this function? h(x) = -2x^2 + 12x + 5 I know the domain is all real numbers, but I'm confused on the range. Thanks.

OpenStudy (anonymous):

h(x) = -2x^2 + 12x + 5 h(x) = -2(x^2-6x-2.5) h(x) = -2(x-3)^2 -11.5 so h is 3, k is -11.5 and a is -2 so range goes from -infinite to -11.5

OpenStudy (anonymous):

Hmm, the answer is saying the range is from -infinity to 23

jimthompson5910 (jim_thompson5910):

h(x) = -2x^2 + 12x + 5 h(x) = -2(x^2 - 6x - 5/2) h(x) = -2(x^2 - 6x + 9 - 9 - 5/2) h(x) = -2((x^2 - 6x + 9) - 9 - 5/2) h(x) = -2((x - 3)^2 - 9 - 5/2) h(x) = -2((x - 3)^2 - 23/2) h(x) = -2(x - 3)^2 - 2(-23/2) h(x) = -2(x - 3)^2 + 23 So the range is (-inf, 23]

OpenStudy (anonymous):

oups, sry

OpenStudy (anonymous):

-infinity to 23 How so i think it is all real numbers.

jimthompson5910 (jim_thompson5910):

If you graph the function, you'll see that it peaks at the point (3,23). So only y values that are equal to 23 or less than it will be part of the range.